# Solving a Real-World Package Delivery Routing Problem Using Quantum Annealers

### v1

@@ Introduction
Figure 6: Step-by-step resolution of D16_P1, consisting of non-priority 15 deliveries and one priority demand (surrounded by a red circle). Also, two non-priority demands belong to the same client (surrounded by a green circle), which are served by the same truck.
Figure 7: Step-by-step resolution of D14_P2, composed non-priority 12 deliveries and two priority demands (surrounded by a red circle). Also, two non-priority demands correspond to the same client (surrounded by a green circle), which must be served separately.
Figure 8: Step-by-step resolution of D21_P2, with 19 non-priority deliveries and two priority ones (surrounded by a red circle).
Performance analysis
[…]
Performance analysis With respect to the Q4RPD performance, it is important to mention that the stage S1 (described in Problem Definition) involved in the pre-selection of the (sub-)route and truck reduces the search space and therefore may affect the global performance of the solver. This means that by attending to the business preferences indicated by Ertransit, the heuristics may hinder or even prevent the optimal solution from being reached. Nevertheless, this stage must be computed in advance to fix the upperbound parameters that complete the mathematical formulation that eventually sets in motion the (sub-)route construction. In addition, this classical processing forces the application of the preferences P1 (i.e., one truck for each complete route) and P2 (i.e., owned trucks have been prioritized to minimize the cost associated with the rental of trucks.). Lastly, and after the conduct of this experimentation, we have found that P3 is not always covered by the currently proposed scheme. Even though this setback only occurs in rare cases, future versions of the scheme will introduce more intelligent resource allocation schemes aiming to meet this preference.
@@
Having said this, the Q4RPD performance in relation to the objective function and constraints for the six instances is summarized in Table 4. Along with the reference to the figure in which the best-found solution by instance is represented, the following information can be found in this table divided into three main blocks:
@@
  • Problem solutions: the amount of complete routes that compose the best solution found by Q4RPD; and the number and type of (sub-)routes of that solution.
@@
  • Validation: in this second section, the 2DH-PDP constraints fulfillment R1 (i.e., respect truck capacities), R2 (i.e., meet time constraints of TP deliveries), and R3 (i.e., respect the length of the working day) is assessed.
@@
  • Objectives of the SRP: concerning the efficiency of the solutions found, and for comparison purposes, a classic TSP has been selected as the baseline algorithm to assess the performance of our solution. Concretely, each route built by Q4RPD is evaluated against a TSP solution obtained through Google OR-Tools. It is important to note that Google OR-Tools only calculates the canonical TSP. Our solver Q4RPD, in turn, is forced to meet the TPs schedule imposed by R1, which causes tweaks in the path that increase the distance but avoids violating this constraint. Specifically, the sum of all the $o_{1}$ objectives calculated by Q4RPD is displayed with a percentage representing the difference with respect to the sum of the costs of the solutions obtained by Google OR-Tools. In addition to that, we also represent in Table 4 if all $o_{2}$ calculated by Q4RPD have been optimized, i.e., if all the built (sub-)routes maximize the number of deliveries attended while meeting restrictions R1 and R3.
@@
Two main conclusions can be drawn from the results. First, Q4RPD has demonstrated to be a promising solver for dealing with 2DH-PDP. Regarding the performance of the solver in terms of constraint fulfillment and its capability to minimize the objective function:
@@
In this paper, a quantum-classical system for solving a real-world logistic problem is presented, named Quantum for Real Package Delivery (Q4RPD). Specifically, this system deals with complex routing problems whose specifications include constraints such as a heterogeneous fleet of vehicles, priority deliveries, and package description and truck capacity according to weight and dimensions. For this reason, the problem has been coined as 2-Dimensional and Heterogeneous Package Delivery with Priorities (2DH-PDP). To demonstrate the applicability, six instances of different sizes and characteristics have been utilized. Several inspiring challenges and opportunities have been identified for future work, graphically summarized and categorized in Fig.11:
@@
    • An extension of the mathematical formulation of 2DH-PDP to include further real-world situations. Some examples of new features would add more realistic and/or complex cost calculations with the cost of the node-to-node trip subject to: a pre-calculated fuel consumption, the likelihood of road congestion…
@@
    • The possibility of reusing the trucks so that they can complete more than one route.
@@
    • The use of complexity reduction mechanisms with the intention of dealing with larger scenarios in an efficient way by reducing the size of the problem while maintaining all the constraints of the problem.
@@
    • The development of more configurable solvers that could automatically handle changes in company preferences or that could easily adapt to other companies requirements. A very interesting avenue would be to grant the solver the ability to transform heuristics, i.e., preferences, into constraints or sub-objectives in the objective function, and the other way around.
Figure 11: A graphical summary of the planned future work.

### v3

@@ Introduction
Having said this, the Q4RPD performance of the objective function and the constraints for the six instances are summarized in Table 4. Along with the reference to the figure in which the best-found solution of the instance is represented, the following information, divided into three main blocks, is provided in this table:
@@
  • Problem solutions: the number of complete routes that compose the best solution found by Q4RPD; and the number and type of (sub-)routes of that solution.
@@
  • Validation: in this second section, the fulfillment of constraints R1 (i.e., truck capacities), R2 (i.e., time constraints of TP deliveries), and R3 (i.e., the length of the working day) is assessed.
  • Objectives of the SRP: concerning the efficiency of the solutions found, and for comparison purposes, a classic TSP has been selected as the baseline algorithm to assess the performance of our solution. Concretely, each route built by Q4RPD is evaluated against a TSP solution obtained from Google OR-Tools, which has been used for this study because of the following reasons: i) it is open-source software; ii) it has been frequently employed for benchmarking purposes [45, 46, 47]; and iii) it has demonstrated a good performance for solving combinatorial optimization problems [48].
@@
    It is important to note that Google OR-Tools only calculates the canonical TSP. Our solver Q4RPD, in turn, is forced to meet the TPs schedule imposed by R1, which causes tweaks in the path that increase the distance but avoids violating this constraint. Specifically, the sum of all the $o_{1}$ objectives calculated by Q4RPD is displayed with a percentage representing the difference with respect to the sum of the costs of the solutions obtained by Google OR-Tools. In addition to that, we also represent in Table 4 if all $o_{2}$ calculated by Q4RPD have been optimized, i.e., if all the generated (sub-)routes maximize the number of deliveries attended while meeting restrictions R1 and R3.
@@
Two main conclusions can be drawn from the results. First, Q4RPD has proved to be a promising solver for dealing with 2DH-PDP. Regarding the performance of the solver in terms of constraint fulfillment and its capability to minimize the objective function:
@@
In this paper, a quantum-classical system for solving a real-world logistic problem is presented, named Quantum for Real Package Delivery (Q4RPD). Specifically, this solving scheme deals with complex routing problems whose specifications include constraints such as a heterogeneous fleet of vehicles, priority deliveries, package description, and truck capacity according to weight and dimensions. For this reason, the problem has been coined as 2-dimensional and Heterogeneous Package Delivery with Priorities (2DH-PDP). To demonstrate the applicability, six instances of different sizes and characteristics have been utilized. Several inspiring challenges and opportunities have been identified for future work, graphically summarized and categorized in Fig.11:
@@
    • An extension of the mathematical formulation of 2DH-PDP to include further real-world situations. Some examples of new features would add more realistic and/or complex cost calculations with the cost of the node-to-node trip subject to pre-calculated fuel consumption, the likelihood of road congestion,
@@
    • The possibility of reusing the trucks to complete multiple routes.
@@
    Figure 11: A graphical summary of the planned future work.
@@
    • The use of complexity reduction mechanisms to deal with larger scenarios efficiently by reducing the problem’s size while maintaining all the constraints of the problem.
@@
    • The development of more configurable solvers that could automatically handle different company preferences or easily adapt to other companies’ requirements. Another interesting avenue would be to grant the solver the ability to transform heuristics, i.e., preferences, into constraints or sub-objectives in the objective function, and the other way around.
    • The implementation of a full-classical algorithm to make a performance comparison with Q4RPD.
@@
Acknowledgements
[…]
Acknowledgements
This work was supported by the Basque Government through HAZITEK program (Q4_Real project, ZE-2022/00033) and through Plan complementario comunicación cuántica (EXP. 2022/01341) (A/20220551). The authors thank Miguel Angel Casla, Ertransit’s Director of Operations, for his assistance throughout this research.
Author contributions statement
[…]
Author contributions statement
All authors conceived the research and defined the problem. All authors conceived the experiments. E.O. and E.V.R design the solver. E.O. developed the code and conducted the experimentation. E.O. and E.V.R. wrote the manuscript. All authors reviewed the manuscript.
Competing interests
[…]
Competing interests
The authors declare no competing interests.